# iPanda: An Intelligent Protocol Testing and Debugging Agent for Conformance Testing

### v1

@@ 3.3.1. Potential issues in LLMs
Although LLMs have demonstrated exceptional performance, their outputs are not entirely reliable. Taking code generation as an example, even when restricting the scope to a certain protocol implementation library, LLMs still exhibit the following issues during the code generation process:
LLM’s limited understanding of code libraries. We analyzed the cutoff dates of training data for several mainstream base LLMs, as shown in Tab. 1. Due to the delay in data acquisition, even the most recent base LLMs are trained on data that is only up to date as of a few months prior. For new implementation libraries, if they are not promptly included in the training dataset, these implementations remain unknown to the LLM. Furthermore, a scarcity of open-source examples for implementation libraries also results in the lack of training data, making it difficult for LLMs to efficiently generate code based on these libraries. This phenomenon will be further demonstrated in subsequent experiments.
@@
Hallucination issues in LLM-generated code. The hallucination problem refers to instances where the LLM generates outputs that appear realistic and credible but are actually incorrect, fabricated, or baseless. This issue stems from the probabilistic prediction mechanisms inherent in statistical learning and is difficult to eliminate completely. In this work, hallucinations in LLM-generated code include calling non-existent classes or attributes from the implementation library, incorrectly using classes or methods in the implementation library, and incorrectly configuring the required parameters.
@@
These issues lead to inefficiencies when blindly relying on LLMs to generate task-specific code based on a given implementation library.
@@
| GPT-4.5-preview (OpenAI, 2025) | 2025/2/27 | 2023/10 |
| GPT-4o(2024-11-20) (Hurst et al., 2024) | 2024/11/20 | 2023/10 |
| Deepseek-R1 (Guo et al., 2025) | 2025/1/20 | unknown |
| Deepseek-V3 (Liu et al., 2024) | 2024/12/26 | unknown |
| Claude-3-7-sonnet (Anthropic, 2025) | 2025/2/19 | 2024/11 |
| Gemini-2.0-flash (Google, 2025) | 2025/2/5 | 2024/1 |
| Qwen2.5-Coder-32B (Hui et al., 2024) | 2024/11/6 | unknown |
3.3.2. codeRAG
Insight: For newly developed protocol implementation libraries, the code details and examples in these libraries naturally contain high-quality, protocol-specific knowledge that LLMs may lack.
LLMs can efficiently learn and utilize new knowledge through in-context learning. This forms the foundation for the effectiveness of Retrieval-Augmented Generation (RAG). To address the issues mentioned above, inspired by the above insight, we design a code-based RAG mechanism named codeRAG, which dynamically guides iPanda in utilizing protocol implementation libraries. As shown in Fig. 3, given a protocol implementation library consisting of source code documents , we first use an embedding model to map these documents into their corresponding embedding representations . To enhance the prompt at the current step , we apply the same embedding model to map it into the same feature space, obtaining its corresponding embedding . The retriever then evaluates the similarity between and other embeddings using the cosine similarity formula:
where the larger the similarity score , the more relevant the document is to the current prompt . By default, we select the four source code documents with the highest similarity scores as reference examples and append them to the prompt for augment. This allows the LLM to leverage few-shot in-context learning to learn how to use the protocol implementation library, leading to more precise code generation.
It is important to note that codeRAG plays a role in both the startup stage (generating code for the first time) and iteration stage (iteratively correcting the code) of iPanda’s code generation process. In the startup stage, the retrieved source code documents help enhance the LLM’s understanding of the protocol implementation. In the subsequent iteration stage, these documents assist in mitigating the LLM’s hallucination issues.
We illustrate the effectiveness of codeRAG with an example. Taking rsocket-py as a case study, when generating code directly from test cases, the LLM may call non-existent methods in the classes due to its limited understanding of rsocket-py. By incorporating the retrieved source code of the relevant class, the LLM can self-correct these bugs. Furthermore, subsequent experiments in Sec. 4.4 also confirm the effectiveness of introducing codeRAG.
As part of codeRAG, the source code of protocol implementation, embedding model, and retriever are integrated into the long-term memory repository of the memory module. At the same time, the long-term memory repository also stores executable code documents generated from previous tests and historical debugging experience summarized by the summarizing module. In addition to retrieving the source code, codeRAG also retrieves these stored knowledge assets, further improving iPanda’s capability in automatic code synthesis and debugging.
3.3.3. Augmented Role Prompting
Prompt engineering is one of the key techniques for improving the quality of LLM’s outputs (White et al., 2023; Giray, 2023; Ekin, 2023). By carefully designing prompts, we can guide the LLM to generate results that better align with expectations. In Sec. 3.3.2, we introduced the optimization technique of using RAG. However, the complete structure of the final prompt used for LLM inference remains unclear. In this sub-section, we present the Augmented Role Prompting method used in iPanda.
In iPanda, we adopt and augment the Role Prompting method to generate prompt templates. Role Prompting explicitly assigns a specific role to the LLM within the prompt, guiding it to perform tasks more effectively, and improving response conformance, accuracy, and professionalism. Generally, a basic Role Prompting structure includes the following components:
Role. It specifies the particular role that the LLM should assume.
Task. It clearly defines the task that the LLM needs to perform. In the startup stage, the task is to generate code, while in the iteration stage, the task is debugging.
Context. It represents the contextual information that the LLM may utilize, i.e., the knowledge retrieved using the codeRAG method from the long-term memory repository, as discussed in Sec. 3.3.2.
In addition to the components mentioned above, the Augmented Role Prompting also introduces instruction to control the LLM’s output. The instruction component primarily includes the following elements:
Zero-shot CoT Prompting. It activates the LLM’s reasoning capabilities to improve the quality of code generation.
Additional guiding statements. They encourage the LLM to actively utilize context knowledge, automate decision-making in code generation, and facilitate debugging.
LLM output requirements. Specifically, the LLM generates outputs in JSON format based on a predefined output template. The JSON output template is structured as follows:
"script1": "print(’Hello, script1’)",
"script2": "print(’Hello, script2’)",
"script3": "print(’Hello, script3’)",
"execution_order": [
["script2", "script3"],
["script1"] ]
, where execution_order controls the execution order of the generated code. The scripts belonging to the same list are executed in parallel.
The Augmented Role Prompting is integrated into the prompt generator of the reasoning module. The complete generated prompt is shown in Fig. 4, where represents the prompt used in the startup stage, while denotes the prompt used in the iteration stage.
3.4. Optimizing code generation
3.4.1. Shortcomings in the startup stage
During the startup stage of iPanda, although the LLM relies on codeRAG and Augmented Role Prompting, effectively reducing the likelihood of hallucinations, we still cannot guarantee that the test code generated by the LLM is both accurate and executable. This issue arises from the inherent limitations of LLMs, which have been thoroughly discussed in Sec. 3.3.1. Therefore, it is necessary to introduce more sophisticated mechanisms to improve the quality of code generation.
3.4.2. Iteration stage: augmented CRITIC
Insight: For code generation, error messages encountered during code execution serve as a naturally accessible and high-quality source of knowledge.
Naturally, the most straightforward way to verify the quality of generated code is to execute it and observe the results. Whether it is the output from a successful execution or a bug report from a failed execution, this information provides additional insights for the LLM. Such feedback can effectively guide the LLM in refining and debugging.
Inspired by this insight, we introduce the iteration stage. The reasoning module’s LLM utilizes an external code execution tool from the execution module to iteratively validate and refine its output through human-like interaction. This process continues until either executable code is generated or the maximum number of retries is reached. This method, known as Self-Correcting with Tool-Interactive Critiquing (CRITIC), was initially proposed in (Alinezhad et al., 2019). Specifically, given the reasoning module’s LLM model and the input test case , the initial code solution is generated by the prompt as . The execution module’s tool then tests this initial solution, producing critiques . For the -th iteration (), the LLM’s output iterates as , where . In the context of code generation, at -th step consists only of the code generated at the current step. Thus, the iterative process does not retain a debugging trajectory of previous code versions. To address this, we augment CRITIC by modifying the LLM’s output at step as:
@@ 3.4.2. Iteration stage: augmented CRITIC
where is a prompt template designed to guide the LLM in debugging. The parameter represents the window size of the short-term memory cache, preventing the input length from exceeding the maximum text length that the LLM can handle (e.g., GPT-4o supports up to 128K tokens (Hurst et al., 2024)). When , the augmented CRITIC degrades to a codeRAG-based CRITIC; when , it considers all historical interaction information. The is the context retrieved using the codeRAG retriever , specifically as:
For code generation, the critiques returned by the execution module play a crucial role in the debugging process, as they identify code errors and provide actionable modification suggestions. These critiques guide the next iteration of code, helping to avoid similar mistakes. Inspired by the human iterative drafting and refinement process, we adopt an iterative validation-correction method. This process continues until the generated code meets specific execution success criteria or the maximum step of iterations is reached. The pseudocode for augmented CRITIC is shown in Alg. 1.
The augmented CRITIC is integrated into the reasoning module, execution module, and memory module. It operates through textual interactions among these modules to achieve iterative refinement. Specifically, the interaction between the reasoning module and the execution module is structured as a JSON object containing the test code and execution sequence, following the JSON template mentioned before. The interaction between the execution module and the memory module also involves JSON objects, which record the execution results or error reports for each generated code. Meanwhile, the interaction between the reasoning module and the memory module is text-based, embedding both the historical JSON strings cached in the short-term memory cache and the code retrieved from the long-term memory repository using the codeRAG.
@@
Remark: During the iteration stage of iPanda, the test code is continuously refined. As the number of iterations increases, the length of the input context grows linearly, which in turn increases the reasoning load on the LLM. Therefore, it is crucial to set an appropriate cache window size to limit the context length. Additionally, as the number of iterations increases, the marginal benefit of code correction diminishes. Hence, selecting an optimal maximum step of iterations is essential. Based on experimental observations, we set the default maximum step to .
3.5. Summarizing bugs
@@
Summarizing code generation experience. Under appropriate prompt guidance, the summarization process generates experience by analyzing the code testing process , where does not exceed the predefined maximum step of LLM iterative reasoning. The experience is stored in the long-term memory repository as valuable knowledge to continuously enhance iPanda’s overall performance. By leveraging the accumulated experience, the augmented CRITIC can refine code corrections more efficiently, ensuring that iPanda continuously learns and improves its ability to handle similar tasks in the future with greater efficiency and accuracy.
@@
Ensuring conformance between protocol and its implementation. With appropriate prompt guidance, the LLM also reviews the test cases and the usage of the protocol implementation library. It evaluates whether the tested protocol implementation adheres to the protocol specifications embedded in the test cases, and generates a corresponding conformance testing report.
Due to the inherent randomness in response generation and the LLM’s imperfect adherence to instructions, the LLM may mistakenly interpret the task of conformance tasting as summarizing bugs encountered during the code generation process. Based on experimental observations, the common types of bugs identified by the LLM in code generation primarily include incorrect parameter passing (e.g., incorrectly binding to any-address) and incorrect method or attribute calls (e.g., some attribute does not exist). To address this issue, a keyword-based filtering method can be applied to the conformance testing reports. This filtering process removes portions of the report that contain keywords related to code generation bugs. The remaining reports are then subjected to manual review and verification to ensure the accuracy of the conformance testing results.
4. Evaluation
@@
In this section, we present our experimental setup, all experimental results and analyses, to demonstrate the effectiveness of iPanda.
4.1. Experimental setup
4.1.1. Experimental platform
Hardware. We deployed the iPanda prototype system on a local server, equipped with a 32-core 13th Gen Intel(R) Core(TM) i9-13900HX CPU and an NVIDIA GeForce RTX 4060 Laptop GPU. To maximize the potential of iPanda, we utilized a cloud-based LLM API for reasoning, meaning that the local device is primarily responsible for managing iPanda’s process control and long-term memory repository. As a result, the demand for CPU and GPU resources is relatively low. The GPU is only required to support the normal operation of the embedding model.
Software. We implemented iPanda using Python. All target protocol implementation libraries, along with their required libraries, were pre-installed in a local virtual environment. For LLM, we used GPT-4o (Hurst et al., 2024), DeepSeek-V3 (Liu et al., 2024), and Qwen2.5-Coder-32B (Hui et al., 2024). These models represent the most advanced base models, differing in architecture, size, and pretraining focus. Notably, we deliberately included a smaller-scale Coder model to examine how a code-specialized model, fine-tuned for coding tasks, adapts to the conformance testing task (Suzgun et al., 2022; Shao et al., 2024). All LLMs were configured with short-term memory window size , , and . However, since the LLM services were accessed via APIs, slight randomness remained in the generated responses. For codeRAG, we utilized OpenAI’s text-embedding-3-large as embedding model (OpenAI, 2024). For simplicity and generality, the testing environment was configured locally, using the local IP address and different ports to simulate network conditions.
4.1.2. Tested protocols and Implementations
We selected the following protocols and their corresponding Python implementation libraries as test subjects:
@@ 4.1.2. Tested protocols and Implementations
CoAP & aiocoap (aiocoap, 2025). The Constrained Application Protocol (CoAP) is a lightweight network communication protocol designed for resource-constrained devices, primarily used for IoT devices. It was standardized in 2014 as RFC 7252 (Shelby et al., 2014). As of March 2025, there are RFC documents related to CoAP (Editor, 2025). We selected aiocoap, the Python implementation of CoAP, as the target for conformance testing. As of March 2025, aiocoap fully or partially supports CoAP-related RFC documents.
@@
RSocket & rsocket-py. RSocket is an application protocol that provides reactive stream semantics over an asynchronous, binary boundary. As of March 2025, RSocket has not yet been formalized as an RFC standard. However, it has a well-established protocol specification (rsocket, 2024) (adhering to RFC 2119) and multiple implementations across various programming languages. We selected rsocket-py, the Python implementation of RSocket, as the target for conformance testing.
The key reasons for selecting these two protocols and their implementation libraries for testing are that they meet important criteria for conformance testing:
The protocols have established mature standards and specifications, but their implementations have only recently developed.
Only partial specifications have been implemented in these protocol libraries.
There is a lack of mature conformance testing tools for these implementations.
It is important to note that LLMs may already contain embedded knowledge about the selected protocols and their implementation libraries. While the RAG is to supplement the LLM with missing knowledge. Therefore, through explicit interaction with the LLM, we inferred the following conclusions. For aiocoap, the LLM possesses extensive built-in knowledge about aiocoap. Consequently, when testing it, we froze the long-term memory repository to prevent unnecessary external influence. For rsocket-py, due to its relatively recent development, and limited open-source examples, the LLM lacks accurate knowledge of it. Therefore, when testing it, we retained the long-term memory repository and incorporated its open-source code. Subsequent experiments confirmed the validity of these operations.
4.1.3. Dataset
@@
Using the keyword-based TCG in Sec. 3.2.2, we introduced two test case sets for CoAP and RSocket:
@@
All test cases underwent professional manual review and selection to ensure their authenticity and validity. Each test case follows a standardized format comprising five components: test case name, test preconditions, test steps, test assertion, and precautions. It is important to note that the core focus of our work is on the accuracy of iPanda’s code generation and the effectiveness of its conformance testing. Therefore, we do not aim for the test case sets to comprehensively cover all protocol specifications. Instead, we ensure the rationality of the test cases and maintain conformance across all experiments by using the same set of test cases.
4.1.4. Baseline

### v2

@@ III-B Automated Program Synthesis
Although LLMs have demonstrated exceptional performance, their outputs are not entirely reliable. For conformance testing, even when restricting the scope to a certain protocol implementation library, LLMs still exhibit the following issues during generating test programs:
LLM’s limited understanding of code libraries. Firstly, LLM’s training data is inherently outdated. As Tab. I shows, even recent models have knowledge cutoffs several months in the past, leaving them unaware of new libraries. Secondly, the scarcity of open-source examples for implementation libraries also results in the lack of training data, hindering effective code generation. This phenomenon will be further demonstrated in subsequent experiments.
@@
Hallucinations in LLM-generated code. LLM hallucinations refer to instances where the LLM generates outputs that appear realistic and credible but are actually incorrect, fabricated, or baseless. This issue, an inherent artifact of the model’s probabilistic nature, is difficult to completely eliminate. In this work, we define code hallucinations as generated code that: (1) calls non-existent library classes or attributes, (2) uses classes or methods incorrectly, or (3) configures required parameters improperly.
@@
These issues lead to inefficiencies when blindly relying on LLMs to generate task-specific programs based on a given implementation library.
@@
| GPT-o3-pro [11] | 2025/6/10 | 2024/6/1 |
| GPT-4o(2024-11-20) [12] | 2024/11/20 | 2023/10 |
| Deepseek-R1-0528 [13] | 2025/5/28 | unknown |
| Deepseek-V3-0324 [14] | 2025/3/25 | unknown |
| Claude-opus-4 [15] | 2025/5/14 | 2025/3 |
| Gemini-2.5-pro [16] | 2025/6/17 | 2025/1 |
| Qwen2.5-Coder-32B [17] | 2024/11/6 | unknown |
1
The statistics were collected on 2025/7/15.
To address these issues, we first introduce code-oriented RAG into the long-term memory repository. We collect all code and examples files related to the implementation. These files naturally contain high-quality, protocol-specific knowledge that LLMs may lack. iPanda then uses OpenAI’s text embedding model to convert these code files into semantic vector representations and builds a local vector database using Chroma [18]. When generating test programs for the test cases, the long-term memory repository first retrieves code files with high similarity to the given test case based on cosine similarity. These retrieved results serve as contexts to help the LLM better understand how to use the implementation.
Moreover, to enable the LLM to generate high-quality test programs, we design a customized CoT method to guide the LLM in the reasoning module to perform step-by-step reasoning [19]. This process simulates how a human developer incrementally develops test subprograms, thereby enhancing the reasoning module’s capability to handle complex test cases.
CoT1: test case understanding. The LLM first decomposes the test case into multiple subtasks based on the test preconditions specified in the case. Specifically, for a test case that involves the collaboration of multiple participants (e.g., clients and servers), the LLM derives a clear and ordered subtask for each participant role and determines the required number of instances for each role. Each role instance comprises a sequence of atomic operations (e.g., connecting to the server rather than connecting and sending data).
CoT2: subprogram generation. Given an instance subtask, the LLM is required to generate the corresponding subprogram, strictly aligning with the subtask operations. The LLM uses the contexts by RAG to understand the implementation and then defines the specific behavior of each role instance in the subprogram. All subprograms are encapsulated as independent executable scripts and stored locally.
CoT3: determining the subprogram execution order. This step orchestrates the independent subprograms into a complete test execution flow. Although each role instance is independent, there are causal dependencies among them. For example, a server cannot accept a connection before a client requests it; a server typically starts before the client and remains in a listening state. Therefore, after all subprograms have been generated, the LLM infers the correct startup sequence based on the test steps defined in the test case.
CoT4: program integration. Once the startup order of the subprograms has been determined, the LLM leverages its in-context learning capability to integrate all subprograms and their corresponding execution order into a single, structured execution blueprint. This blueprint is output as a JSON string, which offers high readability, is easy to parse, and ensures cross-platform compatibility.
III-C Optimizing Test Programs
Automated Test Execution. When conducting automated tests, once the reasoning module generates the integrated program, testers typically need to manually launch all subprograms sequentially. This frequent user intervention is both cumbersome and time-consuming. To address this issue, we develop an automated testing platform within the execution module. This platform automatically parses program texts in JSON format and adaptively launches subprograms in an isolated virtual environment, ensuring that the subprograms are executed safely and under control. If an execution error occurs (e.g. missing attributes and syntax errors), iPanda embeds the execution log into the prompt and initiates its iterative optimization mechanism to debug the program code until the generated program executes successfully. Once the program runs without errors, the LLM evaluates the success of the test by comparing the test assertions defined in the test case with the executed program and its results.
Iterative Optimization Algorithm. Error logs encountered during program execution serve as a naturally accessible and high-quality source of knowledge. They can effectively guide the LLM in refining and debugging. Therefore, iPanda utilizes automated testing platform to iteratively validate and refine its output through human-like interaction [20]. This process continues until either executable program is generated or the maximum number of retries is reached.
Specifically, given the reasoning module’s LLM model and the input test case , the initial program solution is generated by the prompt as . The automated testing platform then tests this initial solution, producing feedback . For the -th iteration (), the LLM’s output iterates as:
@@ III-C Optimizing Test Programs
where is a prompt template designed to guide the LLM in debugging. The parameter represents the window size of the short-term memory cache, preventing the input length from exceeding the maximum text length that the LLM can handle (e.g., GPT-4o supports up to 128K tokens [12]). When , it considers all historical interaction information. The is the context retrieved from long-term memory repository by , specifically as:
This optimization process continues until the generated program meets specific execution success criteria or the maximum step of iterations is reached. The pseudocode for iterative optimization algorithm is shown in Alg. 1.
@@
Remark: During the iteration, the test program is continuously refined. As the number of iterations increases, the length of the input context grows linearly, which in turn increases the reasoning load on the LLM. Therefore, it is crucial to set an appropriate cache window size to limit the context length. Additionally, the more iterations, the smaller the marginal benefit of program correction. Hence, selecting an optimal maximum step of iterations is essential. Based on experimental observations, we set the default maximum step to .
III-D Summarizing Bugs
@@
Summarizing program generation experience. With appropriate prompt guidance, the module generates experience by analyzing the program testing process . The experience is stored in the long-term memory repository as valuable knowledge to continuously enhance iPanda’s overall performance. By leveraging the accumulated experience, the iterative optimization algorithm can refine program corrections more efficiently, ensuring that iPanda continuously learns and improves its ability to handle similar tasks in the future with greater efficiency and accuracy.
@@
Ensuring conformance between protocol and its implementation. With prompts guidance, the module also reviews the test cases and the usage of the protocol implementation. It evaluates whether the tested protocol implementation adheres to the specifications embedded in the test cases, and generates corresponding conformance testing reports.
A primary challenge is distinguishing genuine conformance testing results from incidental code generation errors. Due to inherent randomness and the imperfect adherence to instructions, the LLM may erroneously report on its own code generation errors (e.g., incorrect parameter passing or invalid attribute calls) instead of actual conformance violations. To isolate the true conformance test results, we implement a two-stage filtering process. Firstly, a keyword-based filter automatically removes reports containing terms associated with common generation bugs, including incorrect parameter passing (e.g., incorrectly binding to any-address) and incorrect method or attribute calling (e.g., some attribute does not exist). Subsequently, the remaining reports undergo manual review to verify their relevance and accuracy.
IV Evaluation
@@
In the section, we present our experimental setup, results and analyses, to demonstrate iPanda’s effectiveness.
IV-A Experimental Setup
IV-A1 Experimental platform
We implemented the iPanda using Python and deployed it on our local server. All target protocol implementations, along with their required libraries, were pre-installed in a local virtual environment. For LLM, we used GPT-4o [12], DeepSeek-V3 [21], and Qwen2.5-Coder-32B [17]. These models represent the most advanced base models, differing in architecture, size, and pretraining focus.
All LLMs were configured with short-term memory window size , , and . For RAG, we utilized OpenAI’s text-embedding-3-large as embedding model [18]. For simplicity and generality, the testing environment was configured locally, using the local IP address and different ports to simulate network conditions.
IV-A2 Tested protocols and Implementations
We selected the following protocols and their Python implementations:
@@ IV-A2 Tested protocols and Implementations
CoAP & aiocoap [22]. The Constrained Application Protocol (CoAP) is a lightweight network communication protocol designed for resource-constrained devices, primarily used for IoT devices. It was standardized in 2014 as RFC 7252 [8]. As of March 2025, there are RFC documents related to CoAP [9]. We selected aiocoap, the Python implementation of CoAP, as the target for conformance testing. As of March 2025, aiocoap fully or partially supports CoAP-related RFC documents.
@@
RSocket & rsocket-py. RSocket is an application protocol that provides reactive stream semantics over an asynchronous, binary boundary. As of March 2025, RSocket has not yet been formalized as a RFC standard. However, it has a well-established specification [4] (adhering to RFC 2119) and implementations across various programming languages. We selected rsocket-py, the Python implementation of RSocket, as the target for conformance testing.
The key reason for selecting them is that they meet important criteria for conformance testing: (1) the protocols have established mature specifications, but their implementations have only recently developed; (2) only partial specifications have been implemented in the implementations; (3) there are no mature conformance testing tools for these implementations.
Notably, LLMs possess varying levels of pre-existing knowledge about specific protocols. Therefore, the RAG was adapted based on the LLM’s prior knowledge of each library. A preliminary assessment revealed that the LLM has extensive built-in knowledge of aiocoap, so we deactivated the long-term memory repository to isolate its inherent capabilities. Conversely, for rsocket-py, a more recent library with limited open-source examples, the LLM lacked sufficient knowledge. Therefore, we retained the long-term memory repository and augmented it with the library’s source code to provide the necessary context.
IV-A3 Dataset
@@
Using the keyword-based TCG in Sec. III-A, we introduced two test case sets for CoAP and RSocket:
@@
All test cases underwent professional manual review and selection to ensure their authenticity and validity. Each test case follows a standardized format comprising five components: test case name, test preconditions, test steps, test assertions, and precautions. Notably, the core of our work is on the accuracy of iPanda’s code generation and the effectiveness of its conformance testing. Therefore, rather than pursuing exhaustive specification coverage, our work prioritizes the validity and relevance of individual test cases. To ensure the consistency and comparability of our findings, we employ the same test cases across all experiments.
IV-A4 Baseline